OpenStudy (anonymous):

How many different 4-person teams can be made from a group of 9 players?

OpenStudy (solomonzelman):

Are you familiar with nCr (combinations) ?

OpenStudy (anonymous):

are you saying about 9C4...

OpenStudy (solomonzelman):

Yes !

OpenStudy (solomonzelman):

This calculator would make it easier, I guess. http://www.calculatorsoup.com/calculators/discretemathematics/combinations.php

OpenStudy (anonymous):

hmmm let me try by this too

OpenStudy (solomonzelman):

sure ;)

OpenStudy (anonymous):

9! / (4!(9 - 4)!)... = 9!/4!(5!)... =9*8*7*6*5*4*3*2*1/9*8*7*6*5*4*3*2*1(5*4*3*2*1)... it comes 120 from this.... what i did wrong... please rectify it...

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \sf \frac{9!}{4! (9 - 4) !} }\) \(\LARGE\color{blue}{ \sf \frac{9!}{4!5!} }\) \(\LARGE\color{blue}{ \sf \frac{1\times 2 \times 3\times 4\times 5\times 6\times 7\times 8\times 9}{4!5!} }\) \(\LARGE\color{blue}{ \sf \frac{1\times 2 \times 3\times 4\times 5\times 6\times 7\times 8\times 9}{(1\times 2 \times 3\times 4\times 5)\times(1\times 2 \times 3\times 4)} }\) \(\LARGE\color{blue}{ \sf \frac{6\times 7\times 8\times 9}{1\times 2 \times 3\times 4} }\) \(\LARGE\color{blue}{ \sf \frac{3024}{24} }\) \(\LARGE\color{blue}{ \sf 126 }\)

OpenStudy (anonymous):

ahh bundle of thanks....

OpenStudy (solomonzelman):

You welcome !